II. Given sigma μ=45 and σ, = 6, determine and plot the location of the z scores of the following raw scores: 56, 70,82
III. The scores of SLU College Admission Test are normally distributed,with = 80 and = 6. Find the following:
1. What is the percentile rank of 84?
2. What percentageof scores fall between 64 and 90?
1
Expert's answer
2020-10-22T18:17:46-0400
"Z=\frac{x-\mu}{\sigma}"

Part I)


"\mu = 45, \ \sigma=6"

When X = 56,


"Z=\frac{56-45}{6}= 1.8333"

When X = 70


"Z=\frac{70-45}{6}=4.1667"

When X = 82



"Z=\frac{83-45}{6}=6.1667"




Part II)



"\mu =80,\ \sigma =6"

1.) At 84,



"Z=\frac{84-80}{6}=0.6667"

At Z= 0.6667, probability is 0.74752.


The percentile score 74.752


2. Between 64 and 90



"\frac{64-80}{6} < Z < \frac{90-80}{6}"

"-2.6667<Z<1.6667"

The corresponding probability is


"P(Z<1.6667)-P(Z>-2.6667)"

"0.95221-0.00383=0.94838"

the probability is 0.9484